OpenStudy (anonymous):

solve each system by the addition method 4x+2y=6 4x-5y=-1 put your answer in ordered pair (x,y)

OpenStudy (jdoe0001):

hint: multiply either by -1 \(\large { \begin{array}{rlcll} 4x+2y=6&&\cancel{ 4x}+2y=6\\ 4x-5y=-1 &{\color{red}{ \times -1}}\implies &\cancel{ -4x}+5y=1 \\\hline\\ &&\square?+\square ?=\square ? \end{array} }\)